OpenStudy (anonymous):

Joshua opened a savings account 11 years ago. The interest rate was 2.9% compounded quarterly. If his balance is currently $1,437.34, how much did he originally deposit into the account?

OpenStudy (anonymous):

a=p(1+r/n)^(nt) 1437.34 = p(1+.029/4)^(4*11)